# Theory of polarization-switchable electrical conductivity anisotropy in nonpolar semiconductors

### v1

@@ Abstract
The anisotropic propagation of particles is a fundamental transport phenomenon in solid state physics. As for crystalline semiconductors, the anisotropic charge transport opens novel designing routes for electronic devices, where the electrical manipulation of anisotropic resistance provides essential guarantees. Motivated by the concept of anisotropic magnetoresistance, we develop an original theory on the electrically manipulatable anisotropic electroresistance. We show that when gaining electric polarization, nonpolar semiconductors can showcase anisotropic electrical conductivities between two perpendicular directions and the electrical conductivity anisotropy (ECA) is switchable by flipping the polarization. By symmetry analysis, we identify several point groups hosting the polarization-switchable ECA. These point groups simultaneously enable polarization-reversal induced conductivity change along specific directions, akin to the tunnelling electroresistance in ferroelectric tunnel junctions. First-principles-based conductivity calculations predict that AlP and NaSrP are two good materials having such exotic charge transport. Our theory can motivate the design of intriguing anisotropic electronic devices (e.g., anisotropic memristor and field effect transistor).
@@
Introduction. The anisotropic transport in crystalline materials — for instance, the anisotropic propagation of magnons Cui et al. (2024); Qi et al. (2023), polaritons Luo et al. (2023), electrons or holes Wang et al. (2019a); Wadley et al. (2016); Jungwirth et al. (2016); Cording et al. (2024); Dai et al. (2022); Zeng et al. (2020); Wang et al. (2019b); Liu et al. (2015); Xia et al. (2014); Benítez et al. (2018) — not only is of fundamental interest to solid state physics, but offers extensive development strategies for electronic, optoelectronic and spintronic devices Zhao et al. (2020); Xia et al. (2014); Baltz et al. (2018); Ritzinger and Výborný (2023); Jungwirth et al. (2016); Benítez et al. (2018). Mostly, the design of such devices relies on the electrical or magnetic manipulation of anisotropic transport Qi et al. (2023); Luo et al. (2023); Wang et al. (2019a); Wadley et al. (2016); Jungwirth et al. (2016); Benítez et al. (2018). A profound example is the anisotropic magnetoresistance that occurs in magnets, where longitudinal electrical resistance therein depends on the direction of magnetic order Baltz et al. (2018); Ritzinger and Výborný (2023); Jungwirth et al. (2016); Walter et al. (2014). In ferromagnetic materials, magnetic field manipulates the magnetization direction, and changes the electrical resistance accordingly Baltz et al. (2018); Jungwirth et al. (2016). This gives birth to early-stage magnetic recording devices Baltz et al. (2018); Ritzinger and Výborný (2023), and yields magnetic sensors extensively deployed in automotive, aerospace, and biomedical realms Ritzinger and Výborný (2023).
@@
The anisotropic magnetoresistance is reminiscent of the conceptual anisotropic electroresistance. Previous works reveal the occurrence of tunnelling electroresistance, a counterpart of tunnelling magnetoresistance, in ferroelectric tunnel junctions, with the electroresistance characterized by the electrical resistance change induced by the reversal of electric polarization Jia et al. (2024); Wu et al. (2020); Xi et al. (2017); Wen et al. (2013); Zhuravlev et al. (2005); Ding et al. (2021); Liu et al. (2016); Hernandez-Martin et al. (2020); Klyukin et al. (2018). With regard to anisotropic electroresistance, a particularly interesting research avenue is to discover crystalline semiconductors with polarization-switchable electrical conductivity anisotropy (ECA). Explicitly, it is desired to find nonpolar semiconductors exhibiting $P$-dependent anisotropic conductivities between two orthogonal $\bm{\chi}$ and $\bm{\chi^{\prime}}$ directions [i.e., $\sigma_{\chi,\chi}(P)\neq\sigma_{\chi^{\prime},\chi^{\prime}}(P)$], so that
@@
where $P$ is an electric polarization induced by external electric field. Equation (1) guarantees that the ECA is manipulable and switchable by external electric field, which promises the design of anisotropic electronic devices (e.g., anisotropic memristor and field effect transistor) Zhao et al. (2020); Wang et al. (2019a); Liu et al. (2015). Besides that, Eq. (1) hints a possibility for a polarization-reversal induced conductivity change along $\bm{\chi}$ (or $\bm{\chi^{\prime}}$) direction in crystalline materials  [^1] According to Eq. (1), the $\sigma_{\chi,\chi}(P)-\sigma_{\chi,\chi}(-P)$ conductivity difference along $\bm{\chi}$ direction equals $\sigma_{\chi^{\prime},\chi^{\prime}}(P)+\sigma_{\chi^{\prime},\chi^{\prime}}(-P)-2\sigma_{\chi,\chi}(-P)$, which is not necessarily zero. The similar logic applies to the conductivity difference along $\bm{\chi^{\prime}}$ direction. In next section, we shall discuss this point in details., mimicking the tunnelling electroresistance in ferroelectric tunnel junctions. In such sense, semiconductors with polarization-switchable ECA may serve as alternates for ferroelectric tunnel junctions in device design.
@@
This work aims to provide guidelines for discovering the aforementioned elusive semiconductors. To this end, we develop a polarization-dependent Boltzmann transport theory within effective mass approximation. Then, we perform symmetry analysis for nonpolar crystallographic point groups, derive their symmetry-adapted energy dispersions, and extract the point groups that host polarization-switchable ECA. Simultaneously, our extracted point groups enable the polarization-reversal driven conductivity change. This is well supported by first-principles-based electrical conductivity calculations, which further predict semiconducting AlP and NaSrP as two representative materials with such exotic transport phenomena. Our theory emphasizes the potential of various nonpolar semiconductors (including the celebrated zinc-blende type semiconductors) in designing devices based on anisotropic charge transport.
@@
Theory of polarization-switchable ECA. We start by developing the polarization-dependent Boltzmann transport theory. Under the approximation of constant relaxation time $\tau$, the longitudinal electrical conductivity (at zero Kelvin), contributed by a polarization-dependent $\epsilon(\mathbf{k},P)$ energy dispersion, is given by Tsirkin (2021); Scheidemantel et al. (2003); Železný et al. (2021); Zhao et al. (2024)
@@
where $\delta$, $\mu$, $P$, and $\mathbf{k}\equiv k_{x}\mathbf{x}+k_{y}\mathbf{y}+k_{z}\mathbf{z}$  [^2] We employ the $xyz$ Cartesian coordinate system with $x$, $y$, and $z$ axes being perpendicular to each other. As a convention, we use $\mathbf{x}$, $\mathbf{y}$, and $\mathbf{z}$ to denote the unit vectors along $x$, $y$, and $z$ axes, respectively. denote Dirac’s $\delta$ function, chemical potential, electric polarization, and electronic momentum, respectively. Writing the electrical conductivity as $\sigma_{\chi,\chi}(P,\mu)$ emphasizes its dependence on electric polarization and chemical potential.
@@
To achieve the polarization-switchable ECA between the $\mathbf{x}$ and $\mathbf{y}$ directions, it is required that $\alpha_{0}=\beta_{0}$, $\alpha_{1}\neq\beta_{1}$  [^3] Our theory is established under the first-order approximation with respect to electric polarization $P$. Beyond such an approximation, this statement is not exactly true. For instance, $\alpha_{P}=\alpha_{0}+\alpha_{1}P+\alpha_{3}P^{3}$ and $\beta_{P}=\alpha_{0}+\beta_{1}P+\beta_{3}P^{3}$ may yield non-zero $\sigma_{x,x}(P,\mu)-\sigma_{y,y}(P,\mu)$, when $\alpha_{1}=\beta_{1}$ and $\alpha_{3}\neq\beta_{3}$., and $\alpha_{P}\beta_{P}\gamma_{P}=\alpha_{-P}\beta_{-P}\gamma_{-P}$ [see Eq. (1)]. Similarly, we can derive the conditions for the polarizaiton-switchable ECA between the $\mathbf{y}$ and $\mathbf{z}$ directions and between the $\mathbf{z}$ and $\mathbf{x}$ directions. Within effective mass approximation, we provide more general discussion on polarization-switchable ECA in Section A of the Supplementary Material (SM).
@@
which suggests the asymmetry for resistances between $+P$ and $-P$ states, with this asymmetry depending on the magnitude of electric polarization.
Symmetry analysis. We now provide guidelines to screen nonpolar and nonmagnetic  [^4] As shown in Eqs. (3)–(6), neither magnetism nor spin-orbit interaction is essential for polarization-switchable ECA. Magnetism and spin-orbit interaction are thus neglected in our derivations. semiconductors enabling the polarization-switchable ECA. To this end, we focus on nonpolar crystallographic point groups, perform symmetry analysis, and derive symmetry-adapted energy dispersions for these groups. Depending on polarization $\mathbf{P}\equiv P_{x}\mathbf{x}+P_{y}\mathbf{y}+P_{z}\mathbf{z}$, the formalism of the energy dispersion is
@@
Table 1: The $\Gamma_{1}$ and $\Gamma_{4}$ irreducible representations for the $\bar{4}m2$ point group poi ; Koster et al. (1963); spa . In this table, $\mathfrak{1}$, $\bar{\mathfrak{4}}_{z}^{+}$, and $\bar{\mathfrak{4}}_{z}^{-}$ are identity operation, four-fold rotoinversion along $\mathbf{z}$ direction, and the inverse of $\bar{\mathfrak{4}}_{z}^{+}$, respectively; $\mathfrak{m}_{x}$ ($\mathfrak{m}_{y}$) is the mirror plane perpendicular to $\mathbf{x}$ ($\mathbf{y}$) direction, and $\mathfrak{2}_{xy}$ ($\mathfrak{2}_{x\bar{y}}$) is the two-fold rotation along $\mathbf{x}+\mathbf{y}$ ($\mathbf{x}-\mathbf{y}$) direction.
@@
In the following, we use $\bar{4}m2$ and $\bar{4}2m$ point groups to exemplify the derivations of the $P_{z}$-dependent energy dispersions. Our derivations are based on the group representation theory (see e.g., Refs. Hergert and Geilhufe (2018); El-Batanouny and Wooten (2008)). As shown in Table 1, the $k_{x}^{2}+k_{y}^{2}$ and $k_{z}^{2}$ bases belong to the identity representation $\Gamma_{1}$. This indicates that $(k_{x}^{2}+k_{y}^{2})$ and $k_{z}^{2}$ are invariant with respect to the $\bar{4}m2$ point group, yielding $\tau_{xx}(k_{x}^{2}+k_{y}^{2})$ and $\tau_{zz}k_{z}^{2}$ terms. The bases for the $\Gamma_{4}$ representation are $P_{z}$ and $k_{x}^{2}-k_{y}^{2}$, which are not invariant (e.g., $\bar{\mathfrak{4}}_{z}^{+}:P_{z}\rightarrow-P_{z},k_{x}^{2}-k_{y}^{2}\rightarrow-k_{x}^{2}+k_{y}^{2}$). Since $\Gamma_{4}\otimes\Gamma_{4}=\Gamma_{1}$, we can combine $P_{z}$ with $k_{x}^{2}-k_{y}^{2}$ to generate a new base for the $\Gamma_{1}$ representation [i.e., $\lambda_{zxx}P_{z}(k_{x}^{2}-k_{y}^{2})$]. Overall, the $P_{z}$-dependent energy dispersion associated with $\bar{4}m2$ is $\epsilon(\mathbf{k},P_{z})=(\tau_{xx}+\lambda_{zxx}P_{z})k_{x}^{2}+(\tau_{xx}-\lambda_{zxx}P_{z})k_{y}^{2}+\tau_{zz}k_{z}^{2}$. According to Eq. (5), this suggests the polarization-switchable ECA between the $\mathbf{x}$ and $\mathbf{y}$ directions. Similarly, we can derive the $P_{z}$-dependent energy dispersion for the $\bar{4}2m$ point group as $\epsilon(\mathbf{k},P_{z})=\tau_{xx}(k_{x}^{2}+k_{y}^{2})+\tau_{zz}k_{z}^{2}+\lambda_{zxy}P_{z}k_{x}k_{y}$ [see “The $\bar{4}2m$ point group” in Section B of our SM]. Such an energy dispersion implies the polarization-switchable ECA between the $\mathbf{x}+\mathbf{y}$ and $-\mathbf{x}+\mathbf{y}$ directions. To show this, we define an $x^{\prime}y^{\prime}z$ Cartesian coordinate system whose basis vectors are $\mathbf{x}^{\prime}=(\mathbf{x}+\mathbf{y})/\sqrt{2}$, $\mathbf{y}^{\prime}=(\mathbf{y}+\mathbf{y})/\sqrt{2}$, and $\mathbf{z}$ (see Fig. S1 of the SM). In the $x^{\prime}y^{\prime}z$ coordinate system, the electronic momentum is written as $\mathbf{k}=k_{x^{\prime}}\mathbf{x}^{\prime}+k_{y^{\prime}}\mathbf{y}^{\prime}+k_{z}\mathbf{z}$, and the energy dispersion associated with $\bar{4}2m$ becomes $\epsilon(\mathbf{k},P_{z})=(\tau_{xx}+\lambda_{zxy}P_{z}/2)k_{x^{\prime}}^{2}+(\tau_{xx}-\lambda_{zxy}P_{z}/2)k_{y^{\prime}}^{2}+\tau_{zz}k_{z}^{2}$. Therefore, the $\bar{4}2m$ point group hosts a polarization-switchable ECA between the $\mathbf{x}^{\prime}$ and $\mathbf{y}^{\prime}$ (i.e., $\mathbf{x}+\mathbf{y}$ and $-\mathbf{x}+\mathbf{y}$) directions.
@@
Table 2: Point groups enabling the polarization-switchable ECA, derived up to first order in electric polarization $P_{w}$ ($w=x,y,z$). The entries below each $P_{w}$ are endowed with $(\chi,\chi^{\prime})$ if ECA between the $\bm{\chi}$ and $\bm{\chi^{\prime}}$ directions is enabled by gaining $P_{w}$ and is switchable by reversing $P_{w}$. Otherwise, the entries are endowed with $\cdots$”. Here, $\chi$ or $\chi^{\prime}$ is written as $u$, $uv$, or $\bar{u}v$ ($u,v$ being $x$, $y$, or $z$), where $uv$ and $\bar{u}v$ denote the $\mathbf{u}+\mathbf{v}$ and $-\mathbf{u}+\mathbf{v}$ directions, respectively.
@@
$\bar{4}$cdots $\cdots$ $(x,y)$; $(xy,\bar{x}y)$
$\bar{4}2m$cdots $\cdots$ $(xy,\bar{x}y)$
$\bar{4}m2$cdots $\cdots$ $(x,y)$
$312$ $(xy,\bar{x}y)$cdots $\cdots$
$321$cdots $(xy,\bar{x}y)$cdots
$\bar{6}m2$ $(xy,\bar{x}y)$cdots $\cdots$
$\bar{6}2m$cdots $(xy,\bar{x}y)$cdots
@@
$\bar{6}$ $(x,y)$; $(xy,\bar{x}y)$ $(x,y)$; $(xy,\bar{x}y)$cdots
$\bar{6}m2$cdots $(x,y)$cdots
$\bar{6}2m$ $(x,y)$cdots $\cdots$

### v2

# Theory of polarization-switchable electrical conductivity anisotropy in piezoelectric and ferroelectric semiconductors

@@ Abstract
The anisotropic propagation of particles is a fundamental transport phenomenon in solid state physics. As for crystalline semiconductors, the anisotropic charge transport opens novel designing routes for electronic devices, where the electrical or magnetic manipulation of anisotropic resistance provides essential guarantees. Motivated by the concept of anisotropic magnetoresistance, we develop an original theory on the electrically manipulatable anisotropic electroresistance. We show that piezoelectrics and ferroelectrics may showcase polarization-dependent anisotropic electrical conductivities between two perpendicular directions and the electrical conductivity anisotropy (ECA) is switchable by flipping the polarization. By symmetry analysis, we identify several point groups hosting the polarization-switchable ECA. These point groups simultaneously enable polarization-reversal induced conductivity change along specific directions, akin to the tunnelling electroresistance in ferroelectric tunnel junctions. First-principles-based conductivity calculations predict that piezoelectric AlP and ferroelectric KH2PO4 are two good semiconductors having such exotic charge transport. Our theory can motivate the design of intriguing anisotropic electronic devices (e.g., anisotropic memristor and field effect transistor).
I I. Introduction
@@
The anisotropic transport in crystalline materials — for instance, the anisotropic propagation of magnons [1, 2], polaritons [3], electrons or holes [4, 5, 6, 7, 8, 9, 10, 11, 12, 13] — not only is of fundamental interest to solid state physics, but offers extensive development strategies for electronic, optoelectronic and spintronic devices [14, 12, 15, 16, 6, 13]. Mostly, the design of such devices relies on the electrical or magnetic manipulation of anisotropic transport [2, 3, 4, 5, 6, 13]. A profound example is the anisotropic magnetoresistance that occurs in magnets, where longitudinal electrical resistance therein depends on the direction of magnetic order [15, 16, 6, 17]. In ferromagnetic materials, magnetic field manipulates the magnetization direction, and changes the electrical resistance accordingly [15, 6]. This gives birth to early-stage magnetic recording devices [15, 16], and yields magnetic sensors extensively deployed in automotive, aerospace, and biomedical realms [16].
@@
The anisotropic magnetoresistance is reminiscent of the conceptual anisotropic electroresistance. Previous works reveal the occurrence of tunnelling electroresistance, a counterpart of tunnelling magnetoresistance, in ferroelectric tunnel junctions, with the electroresistance characterized by the electrical resistance change induced by the reversal of electric polarization [18, 19, 20, 21, 22, 23, 24, 25, 26]. With regard to anisotropic electroresistance, a particularly interesting research avenue is to discover crystalline semiconductors with polarization-switchable electrical conductivity anisotropy (ECA). Explicitly, it is desired to find semiconductors exhibiting $P$-dependent anisotropic conductivities between two orthogonal $\bm{\chi}$ and $\bm{\chi^{\prime}}$ directions [i.e., $\sigma_{\chi,\chi}(P)\neq\sigma_{\chi^{\prime},\chi^{\prime}}(P)$], so that
@@
where $P$ is an electric polarization induced by an electric field or spontaneously occurred. Equation (1) guarantees that the ECA is manipulable and switchable by external electric field, which promises the design of anisotropic electronic devices (e.g., anisotropic memristor and field effect transistor) [14, 4, 11]. Besides that, Eq. (1) hints a possibility for a polarization-reversal induced conductivity change along $\bm{\chi}$ (or $\bm{\chi^{\prime}}$) direction in crystalline materials  [^1] According to Eq. (1), the $\sigma_{\chi,\chi}(P)-\sigma_{\chi,\chi}(-P)$ conductivity difference along $\bm{\chi}$ direction equals $\sigma_{\chi^{\prime},\chi^{\prime}}(P)+\sigma_{\chi^{\prime},\chi^{\prime}}(-P)-2\sigma_{\chi,\chi}(-P)$, which is not necessarily zero. The similar logic applies to the conductivity difference along $\bm{\chi^{\prime}}$ direction. In next section, we shall discuss this point in details., mimicking the tunnelling electroresistance in ferroelectric tunnel junctions. That is, semiconductors with polarization-switchable ECA may serve as alternates for ferroelectric tunnel junctions in device design.
@@
The present work aims to provide guidelines for discovering the aforementioned elusive semiconductors. For this, we develop a polarization-dependent Boltzmann transport theory within effective mass approximation. Then, we perform symmetry analysis for crystallographic point groups, and extract the members that host polarization-switchable ECA. These point groups are associated with piezoelectric and ferroelectric materials. Simultaneously, our extracted point groups enable polarization-reversal driven conductivity change. This is further corroborated by first-principles-based electrical conductivity calculations, which helps to identify piezoelectric AlP and ferroelectric KH2PO4 as two representative semiconductors with such exotic transport phenomena. Our theory emphasizes the potential of various piezoelectric and ferroelectric semiconductors (including the celebrated zinc-blende type semiconductors) in designing devices based on anisotropic charge transport.
II II. Methods
We perform symmetry analyses, sketch drawings, and data plots with the help of Bilbao Crystallographic Server [28, 29, 30], SeeK-path [31, 32], Mathematica [33], Matplotlib [34], Pymatgen [35], and VESTA [36]. For numerical simulations, we employ the VASP [37, 38] code (based on PAW potentials [39]) to perform first-principles calculations and the Boltztrap2 [40] code to carry out charge transport calculations. We use the PBEsol exchange-correlation functional [41] without involving spin-orbit interactions, and solve the following electronic configurations: $1s$ for H, $(2p,3s)$ for Na, $(3s,3p,4s)$ for K, $(4s,4p,5s)$ for Sr, $(3s,3p)$ for Al, $(2s,2p)$ for O and $(3s,3p)$ for P. We select the kinetic energy cutoff values of 450 eV, 550 eV and 450 eV for AlP, KH2PO4, and NaSrP, respectively. The crystal structures of $F\bar{4}3m$ AlP, $Fdd2$ KH2PO4 and $P\bar{6}2m$ NaSrP are schematized in Fig. 1 and Fig.S2 of the Supplementary Material (SM)  [^2] See Supplementary Material at a link which contains general remarks on polarization-switchable ECA, symmetry analysis on ECA, and some numerical results., where the SM also contains Refs. [43, 44]. In particular, we work with the super cell of AlP whose cell volume is twice that of its unit cell [see Fig. 1(a)]. The $k$-point meshes are $8\times 8\times 12$ for AlP, $6\times 6\times 10$ for KH2PO4, and $8\times 8\times 14$ for NaSrP for sampling the Brillouin zone, during the structural relaxations or band structure calculations. We carry out structural relaxations with the force convergence criterion of 1 meV/Å, where the approach developed by Bellaiche and Fu [45] enables the calculations of crystal structures under finite electric fields [46, 47]. Starting from the relaxed crystal structures of AlP, KH2PO4 and NaSrP, we do self consistent field calculations to prepare the input files for the transport calculations; For accuracy, the $k$-point meshes are further increased to $60\times 60\times 90$ for AlP, $40\times 40\times 60$ for KH2PO4 and $60\times 60\times 104$ for NaSrP. The transport calculations are performed with a thermal smearing of 300 Kelvin, which yields non-zero conductivity values at the valence band maximum (VBM) or conduction band minimum (CBM). During the transport calculations, the crystal structures at various chemical potentials are kept fixed at their undoped structures. Even though free carriers create electrostatic screening, the polar distortions of a broad range of typical ferroelectrics well survive free carriers when the carrier concentration does not exceed $10^{20}$ cm-3 [48] — an upper limit associated with chemical doping [49]. This implies that a free carrier concentration of up to $10^{20}$ cm-3 is insufficient to prevent the intrinsic or induced polar distortions in materials.
III III. Results and discussion
III.1 A. Theory of polarization-switchable ECA
@@
We start by developing the polarization-dependent Boltzmann transport theory. Under the approximation of constant relaxation time $\tau$, the longitudinal electrical conductivity (at zero Kelvin), contributed by a polarization-dependent $\epsilon(\mathbf{k},P)$ energy dispersion, is given by [50, 51, 52, 53]
@@
where $\delta$, $\mu$, $P$, and $\mathbf{k}\equiv k_{x}\mathbf{x}+k_{y}\mathbf{y}+k_{z}\mathbf{z}$  [^3] We employ the xyzxyz Cartesian coordinate system with $x$, $y$, and $z$ axes being perpendicular to each other. As a convention, we use $\mathbf{x}$, $\mathbf{y}$, and $\mathbf{z}$ to denote the unit vectors along $x$, $y$, and $z$ axes, respectively. denote Dirac’s $\delta$ function, chemical potential, electric polarization, and electronic momentum, respectively. Writing the electrical conductivity as $\sigma_{\chi,\chi}(P,\mu)$ emphasizes its dependence on electric polarization and chemical potential.
@@
To achieve the polarization-switchable ECA between the $\mathbf{x}$ and $\mathbf{y}$ directions, it is required that $\alpha_{0}=\beta_{0}$, $\alpha_{1}\neq\beta_{1}$  [^4] Our theory is established under the first-order approximation with respect to electric polarization $P$. Beyond such an approximation, this statement is not exactly true. For instance, $\alpha_{P}=\alpha_{0}+\alpha_{1}P+\alpha_{3}P^{3}$ and $\beta_{P}=\alpha_{0}+\beta_{1}P+\beta_{3}P^{3}$ may yield non-zero $\sigma_{x,x}(P,\mu)-\sigma_{y,y}(P,\mu)$, when $\alpha_{1}=\beta_{1}$ and $\alpha_{3}\neq\beta_{3}$., and $\alpha_{P}\beta_{P}\gamma_{P}=\alpha_{-P}\beta_{-P}\gamma_{-P}$ [see Eq. (1)]. Similarly, we can derive the conditions for the polarizaiton-switchable ECA between the $\mathbf{y}$ and $\mathbf{z}$ directions and between the $\mathbf{z}$ and $\mathbf{x}$ directions. Within effective mass approximation, we provide more general discussion on polarization-switchable ECA in Section A of the SM.
@@
which suggests the asymmetry for resistances between +P+P and P-P states, with this asymmetry depending on the magnitude of electric polarization.
III.2 B. Symmetry analysis
We now provide guidelines to screen nonmagnetic  [^5] As shown in Eqs. (3)–(6), neither magnetism nor spin-orbit interaction is essential for polarization-switchable ECA. Magnetism and spin-orbit interaction are thus neglected in our derivations. semiconductors enabling polarization-switchable ECA. To this end, we perform symmetry analysis for crystallographic point groups. We first work with nonpolar point groups and extend our discussion to polar point groups at a later stage. Depending on polarization $\mathbf{P}\equiv P_{x}\mathbf{x}+P_{y}\mathbf{y}+P_{z}\mathbf{z}$, the formalism of the energy dispersion is written as
@@
Table 1: The $\Gamma_{1}$ and $\Gamma_{4}$ irreducible representations for the $\bar{4}m2$ point group [57, 58, 59]. In this table, $\mathfrak{1}$, $\bar{\mathfrak{4}}_{z}^{+}$, and $\bar{\mathfrak{4}}_{z}^{-}$ are identity operation, four-fold rotoinversion along $\mathbf{z}$ direction, and the inverse of $\bar{\mathfrak{4}}_{z}^{+}$, respectively; $\mathfrak{m}_{x}$ ($\mathfrak{m}_{y}$) is the mirror plane perpendicular to $\mathbf{x}$ ($\mathbf{y}$) direction, and $\mathfrak{2}_{xy}$ ($\mathfrak{2}_{x\bar{y}}$) is the two-fold rotation along $\mathbf{x}+\mathbf{y}$ ($\mathbf{x}-\mathbf{y}$) direction.
@@
In the following, we use $\bar{4}m2$ and $\bar{4}2m$ point groups to exemplify the derivations of the $P_{z}$-dependent energy dispersions. Our derivations are based on the group representation theory (see e.g., Refs. [60, 61]). As shown in Table 1, the $k_{x}^{2}+k_{y}^{2}$ and $k_{z}^{2}$ bases belong to the identity representation $\Gamma_{1}$. This indicates that $(k_{x}^{2}+k_{y}^{2})$ and $k_{z}^{2}$ are invariant with respect to the $\bar{4}m2$ point group, yielding $\tau_{xx}(k_{x}^{2}+k_{y}^{2})$ and $\tau_{zz}k_{z}^{2}$ terms. The bases for the $\Gamma_{4}$ representation are $P_{z}$ and $k_{x}^{2}-k_{y}^{2}$, which are not invariant (e.g., $\bar{\mathfrak{4}}_{z}^{+}:P_{z}\rightarrow-P_{z},k_{x}^{2}-k_{y}^{2}\rightarrow-k_{x}^{2}+k_{y}^{2}$). Since $\Gamma_{4}\otimes\Gamma_{4}=\Gamma_{1}$, we can combine $P_{z}$ with $k_{x}^{2}-k_{y}^{2}$ to generate a new base for the $\Gamma_{1}$ representation [i.e., $\lambda_{zxx}P_{z}(k_{x}^{2}-k_{y}^{2})$]. Overall, the $P_{z}$-dependent energy dispersion associated with $\bar{4}m2$ is $\epsilon(\mathbf{k},P_{z})=(\tau_{xx}+\lambda_{zxx}P_{z})k_{x}^{2}+(\tau_{xx}-\lambda_{zxx}P_{z})k_{y}^{2}+\tau_{zz}k_{z}^{2}$. According to Eq. (5), this suggests the polarization-switchable ECA between the $\mathbf{x}$ and $\mathbf{y}$ directions. Similarly, we can derive the $P_{z}$-dependent energy dispersion for the $\bar{4}2m$ point group as $\epsilon(\mathbf{k},P_{z})=\tau_{xx}(k_{x}^{2}+k_{y}^{2})+\tau_{zz}k_{z}^{2}+\lambda_{zxy}P_{z}k_{x}k_{y}$ [see “The $\bar{4}2m$ point group” in Section B of our SM]. Such an energy dispersion implies the polarization-switchable ECA between the $\mathbf{x}+\mathbf{y}$ and $\mathbf{x}+\mathbf{y}$ directions. To show this, we define an $x^{\prime}y^{\prime}z$ Cartesian coordinate system whose basis vectors are $\mathbf{x}^{\prime}=(\mathbf{x}+\mathbf{y})/\sqrt{2}$, $\mathbf{y}^{\prime}=(-\mathbf{x}+\mathbf{y})/\sqrt{2}$, and $\mathbf{z}$ (see Fig. S1 of the SM). In $x^{\prime}y^{\prime}z$ coordinate system, the electronic momentum is written as $\mathbf{k}=k_{x^{\prime}}\mathbf{x}^{\prime}+k_{y^{\prime}}\mathbf{y}^{\prime}+k_{z}\mathbf{z}$, and the energy dispersion associated with $\bar{4}2m$ becomes $\epsilon(\mathbf{k},P_{z})=(\tau_{xx}+\lambda_{zxy}P_{z}/2)k_{x^{\prime}}^{2}+(\tau_{xx}-\lambda_{zxy}P_{z}/2)k_{y^{\prime}}^{2}+\tau_{zz}k_{z}^{2}$. Therefore, the $\bar{4}2m$ point group hosts a polarization-switchable ECA between $\mathbf{x}^{\prime}$ and $\mathbf{y}^{\prime}$ (i.e., $\mathbf{x}+\mathbf{y}$ and $\mathbf{x}+\mathbf{y}$).
@@
Table 2: Point groups enabling the polarization-switchable ECA, derived up to first order in electric polarization $P_{w}$ ($w=x,y,z$). The entries below each $P_{w}$ are endowed with $(\chi,\chi^{\prime})$ if ECA between the $\bm{\chi}$ and $\bm{\chi^{\prime}}$ directions is enabled by gaining $P_{w}$ and is switchable by reversing $P_{w}$. Otherwise, the entries are endowed with $\cdots$”. Here, $\chi$ or $\chi^{\prime}$ is written as $u$, uvuv, or $\bar{u}v$ ($u,v$ being $x$, $y$, or $z$), where $uv$ and $\bar{u}v$ denote the $\mathbf{u}+\mathbf{v}$ and $\mathbf{u}+\mathbf{v}$ directions, respectively.
@@
$\bar{4}$ $\cdots$ $\cdots$ $(x,y)$; $(xy,\bar{x}y)$
$\bar{4}2m$ $\cdots$ $\cdots$ $(xy,\bar{x}y)$
$\bar{4}m2$ $\cdots$ $\cdots$ $(x,y)$
$312$ $(xy,\bar{x}y)$ $\cdots$ $\cdots$
$321$ $\cdots$ $(xy,\bar{x}y)$ $\cdots$
$\bar{6}m2$ $(xy,\bar{x}y)$ $\cdots$ $\cdots$
$\bar{6}2m$ $\cdots$ $(xy,\bar{x}y)$ $\cdots$
@@
$\bar{6}$ $(x,y)$; $(xy,\bar{x}y)$ $(x,y)$; $(xy,\bar{x}y)$ $\cdots$
$\bar{6}m2$ $\cdots$ $(x,y)$ $\cdots$
$\bar{6}2m$ $(x,y)$ $\cdots$ $\cdots$